OpenStudy (anonymous):

The asteroid Hygiea orbits the Sun like the other planets. Its period is about 2030 days. What is its mean distance from the Sun?

OpenStudy (wolf1728):

This requires Kepler's Third Law. T^2 = r^3 When dealing with orbits around the Sun, "T" should be stated in years and "r" should be stated in Earth Sun distances (astronomical units). 2,030 days divided by 365.25 = 5.5578 years r^3 = 5.5578^2 r^3 = 30.889 r = cube root (30.889) r = 3.1376 astronomical units which equals 291,660,000 miles = 469,380,000 kilometers And to be very sure of that answer here is a Kepler's Third Law Calculator: http://www.1728.org/kepler3a.htm